Journalist Report – January 30th

30th January, 2019.
Sol 2:

Good Morning from Mars!

The Morning had never been so fascinating when you get an opportunity to observe a completely new horizon from a tiny window. It looks like the sun has spread his arms to hug the mountains and the mountains blush and become red! This is the Crew Journalist and Health and Safety Officer(HSO) Sonal Baberwal reporting live from the lower deck of the habitat observing the beauty of nature.

Preparing and serving powdered food in breakfast, are we really on Mars or on the preparation to be on Mars? What do you think does the powdered food taste like? Well, let me tell you there is an advancement. From the time where there was no flavour in the paste food served to astronauts till the time powdered food had got a great taste.

By the way, did I mention about preparation? Yes, we are preparing for bringing new life here on Mars. The tiny seeds are getting prepared for germination and see the Martian world around them. Looking forward to seeing the GreenHab officer and the crew trying some experiments with the soil in hope for the leaves of the plant to rise over this soil.

And the most awaiting moment comes here. The very first EVA experience, the very first experience to wear the spacesuit. I feel the words are becoming less to describe the emotions of having this experience and getting dressed. It feels more exciting than a bride getting dressed for her wedding!

We had 2 EVAs. Wait a minute, we had only one EVA planned, so where did a sudden EVA come from? Don’t you think it’s like a fairytale that you have been invited for an EVA on a special authorization by Mission Support for a special guest and photographer Mr. Jad Davenport. The first one was performed by Kunal (Space Bot) and Avishek (carbon flight) to an opportunity to participate in the photography session. This was followed by the EVA by Patil (cosmos) and Sonal (Mars bag) to collect soil samples.

The day went busy but awesome and with a lot of exciting activities. Looking forward to the new day with new hope! Stay connected with the Crew204 and we invite you to be the part of the mission towards the destination!

Sonal Baberwal, Crew 204 Journalist

Sol Summary – January 30th

Crew 204 Sol Summary Report 30-JAN-2019

Sol: 2

Summary Title: The Hush and the Rush

Author’s name: Kunal Naik (Space_bot)

Mission Status: Nominal

Sol Activity Summary: The most beautiful thing on Mars is the dawn.
The crew woke up early by 6:30 am and started the day with morning
activities and Martian breakfast. We have requested prior 2 EVA’s one
in the morning and later by evening. Our 1st EVA was delayed and
rescheduled due to limited water tank issue which was resolved at the
earliest. The crew members initiated their project work along with
their Hab responsibilities. Post Lunch we were surprised and glad to
meet the Nat Geo Photographer Jad Davenport who came to visit the
facility for reporting and clicking wonderful photos with the crew. We
had our 1st EVA by 1400 and later 2nd by 1500, we were followed by the
enthusiastic photographer to capture our EVA activities as well as our
moments via his personal camera and his drone. It was a hush and rush day but went well with no setbacks.

Look Ahead Plan: We are going to follow up on our EVA plans as per
today’s investigation and exploration. We got some soil samples during our 2nd EVA which we need to mesh and set some composition parameters for Martian 3D Printer. We have planned to work on fixing Wifi issues which are a small problem for now.

Anomalies in work: None Encountered

Weather: Sunny

Crew Physical Status: Nominal

EVA: Both the EVA’s same location (518750, 4251250)

Reports to be filed: 2 EVA’s done along with Nat Geo Photographer at
the same location due to short time and photo session. Same rover
Spirit was used for both EVA’s as the location was quite near.

Support Requested: None

Operations Report – January 30th

Crew 204 Operations Report 30-Jan-2019

SOL: 2

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): started at 0730 in the Morning, stopped by 1830.

Solar— SOC % Turned on (29th January) 85%, Turned on (30th January) 74%

Diesel Reading – 1/2 Tank

Propane Reading – 73%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – 90%, 450 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – 278.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01406324 units

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Spirit rover used: Used for 1st EVA

Hours:71.6 hours
Beginning charge: 100%
Ending charge: 88%
Currently charging: Yes

Spirit rover used: Used for 2nd EVA (Due to limited time and Natgeo Photos)
Hours:71.8 hours
Beginning charge: 100%
Ending charge: 79%
Currently charging: Yes

Curiosity rover used: Not Used
Hours: Not Used
Beginning charge: Not Used
Ending charge: Not Used
Currently charging: Not Used

Opportunity rover used: Brakes not functional
Hours: 46.2 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? For water

Crew Car used and why, where? Off-Campus

General notes and comments: We have received 2 packages.

Summary of internet: N/A

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Received a box package of
Measuring Cylinder in the science room which was broken by crew 203 which are now in the HAB with the crew 204.

Summary of RAMM operations: Received a box package of MDRS Sim Suit Recharger which is now in the HAB with the crew 204.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: Nothing to report.

EVA Report – January 30th

Crew 204 EVA #2 Report 30 JAN 2019

EVA # 2 (Afternoon)

Author: Avishek Ghosh.

Purpose of EVA: A general exploration and soil sample collection as per planned EVA. The crew also had an opportunity to participate in a photography session invited by a special guest Mr. Jad Davenport.

Start time: 13:00 hrs.

End time: 15:45 hrs.

Narrative: It was a planned EVA. A wonderful opportunity to exhibit the MDRS crew’s preparation for the simulation. The rest two crews also had an opportunity to participate in a photography session by Mr. Jad Davenport.

Destination: A low mountain region just after Pooh’s corner and before the stream bed

EVA Report – January 30th

Crew 204 EVA #1 Report 30 JAN 2019

EVA # 1 (Afternoon)

Author: Avishek Ghosh.

Purpose of EVA: A general exploration. The crew 204 had an opportunity to participate in a photography session invited by a special guest Mr. Jad Davenport.

Start time: 14:00 hrs.

End time: 15:00 hrs.

Narrative: Although an unplanned EVA but, it was a wonderful opportunity to exhibit the MDRS crew’s preparation for the simulation. And also crew 204 feels honoured to participate in a session invited by personnel associated with a prestigious organization like National Geographic.

Destination: A low mountain region just after Pooh’s corner and before the Stream bed connector on Cow dung road, 0110.

Coordinates: X-518750; Y-4251250

Participants: Avishek Ghosh & Kunal Naik.

Roads and routes per MDRS map: Pooh’s Corner.

Mode of travel: Rover Spirit.

Operations Report – January 30th

Crew 204 Operations Report 30-Jan-2019

SOL: 2

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): started at 0730 in the Morning, stopped by 1830.

Solar— SOC % Turned on (29th January) 85%, Turned on (30th January) 74%

Diesel Reading – 1/2 Tank

Propane Reading – 73%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – 90%, 450 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab –200.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01406324 units

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Spirit rover used: Used for 1st EVA
Hours:71.6 hours
Beginning charge: 100%
Ending charge: 88%
Currently charging: Yes

Spirit rover used: Used for 2nd EVA (Due to limited time and Natgeo Photos)
Hours:71.8 hours
Beginning charge: 100%
Ending charge: 79%
Currently charging: Yes

Curiosity rover used: Not Used
Hours: Not Used
Beginning charge: Not Used
Ending charge: Not Used
Currently charging: Not Used

Opportunity rover used: Brakes not functional
Hours: 46.2 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? For water

Crew Car used and why, where? Off-Campus

General notes and comments: We have received 2 packages.

Summary of internet: N/A

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Received a box package of
Measuring Cylinder in the science room which was broken by crew 203 which are now in the HAB with the crew 204.

Summary of RAMM operations: Received a box package of MDRS Sim Suit Recharger which is now in the HAB with the crew 204.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: Nothing to report.

Greenhab Report – January 30th

Crew 204 Green Hab Report – 30-JAN-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.2 °C

High: 28.1°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 9.5 gallons

Water in Blue Tank: ~ 280.5 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 07:40 AM 21.2° C / 42 %
2) 10:53 AM 25.8° C / 22 %
3) 15:00 PM 24.0° C / 22 %
4) 18:22 PM 21.6° C / 46%

Changes to crops: None

Narrative: It was a learning day and while performing EVAs, the crew
has shown a good collaboration to maintain adequate temperature and watering schedules in GreenHab. Mixed Seeds have been kept for soaking and sprouting, which include mat bean, kidney bean, fenugreek, and sesame.

Harvest: None

Support/supplies needed: None